The FBI wants to determine the effectiveness of their 10 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught.

Step 2 of 2 :

Suppose a sample of 465 suspected criminals is drawn. Of these people, 111 were captured. Using the data, construct the 85% confidence interval for the population proportion of people who are captured after appearing on the 10 Most Wanted list. Round your answers to three decimal places.
Answer #1

Step 2: z value for 85% CI is 1.44 as P(-1.44<z<1.44)=0.85

P = 465 = 0.239

Margin of Error is E = 2* 1.44 * 1 0.239 * 0.761 465 = 0.028 n

Hence CI is CI = p+E=0.239 + 0.028 = (0.211, 0.267)